### v2

# SwiLTra-Bench: The Swiss Legal Translation Benchmark

Abstract
In Switzerland legal translation is uniquely important due to the country’s four official languages and requirements for multilingual legal documentation. However, this process traditionally relies on professionals who must be both legal experts and skilled translators—creating bottlenecks and impacting effective access to justice. To address this challenge, we introduce SwiLTra-Bench, a comprehensive multilingual benchmark of over 180K aligned Swiss legal translation pairs comprising laws, headnotes, and press releases across all Swiss languages along with English, designed to evaluate LLM-based translation systems. Our systematic evaluation reveals that frontier models achieve superior translation performance across all document types, while specialized translation systems excel specifically in laws but under-perform in headnotes. Through rigorous testing and human expert validation, we demonstrate that while fine-tuning open SLMs significantly improves their translation quality, they still lag behind the best zero-shot prompted frontier models such as Claude-3.5-Sonnet. Additionally, we present SwiLTra-Judge, a specialized LLM evaluation system that aligns best with human expert assessments.
listing[Listing][List of Listings]
SwiLTra-Bench: The Swiss Legal Translation Benchmark
Joel Niklaus Jakob Merane Luka Nenadic Sina Ahmadi Yingqiang Gao Cyrill A. H. Chevalley Claude Humbel Christophe Gösken Lorenzo Tanzi Thomas Lüthi Stefan Palombo Spencer Poff Boling Yang Nan Wu Matthew Guillod Robin Mamié Daniel Brunner Julio Pereyra Niko Grupen Harvey ETH Zurich Swiss Federal Supreme Court University of Zurich University of Basel University of Geneva University of Lausanne Canton of Solothurn Max Planck Institute for Research on Collective Goods Correspondence: joel@niklaus.ai
- FSCS
- Federal Supreme Court of Switzerland
- SCI
- Supreme Court of India
- ECHR
- European Convention of Human Rights
- ECtHR
- European Court of Human Rights
- SCOTUS
- Supreme Court of the United States
- SPC
- Supreme People’s Court of China
- SJP
- Swiss-Judgment-Prediction
- ASO
- Almost Stochastic Order
- ILDC
- Indian Legal Documents Corpus
- US
- United States
- EU
- European Union
- NLP
- Natural Language Processing
- ML
- Machine Learning
- LJP
- Legal Judgment Prediction
- SJP
- Swiss-Judgment-Prediction
- PJP
- Plea Judgment Prediction
- BERT
- Bidirectional Encoder Representations from Transformers
- LSTM
- Long Short-Term Memory
- GRU
- Gated Recurrent Unit
- BiLSTM
- Bidirectional Long Short-Term Memory
- CNN
- Convolutional Neural Networks
- PLM
- pre-trained Language Model
- LM
- Language Model
- RTD
- Replaced Token Detection
- CLT
- Cross-Lingual Transfer
- HRL
- high resource language
- LRL
- low resource language
- POS
- Part-of-Speech
- SLTC
- Single Label Text Classification
- MLTC
- Multi Label Text Classification
- TC
- Text Classification
- NLU
- Natural Language Understanding
- IR
- Information Retrieval
- NER
- Named Entity Recognition
- NLU
- Natural Language Understanding
- QA
- Question Answering
- NLI
- Natural Language Inference
- GNB
- Gaussian Naive Bayes
- DT
- Decision Tree
- SVM
- Support-Vector Machine
- RF
- Random Forest
- XGBoost
- eXtreme Gradient Boosting
- MLIR
- Multilingual Information Retrieval
- IR
- Information Retrieval
- NDCG
- Normalized Discounted Cumulative Gain
- LD
- Leading Decision
- FSCD
- Federal Supreme Court Decisons
- SFCS
- Swiss Federal Supreme Court
- LLM
- Large Language Model
- LM
- Language Model
- CVG
- Court View Generation
- JP
- Judgment Prediction
- LAP
- Law Area Prediction
- SLAP
- Sub Law Area Prediction
- CP
- Criticality Prediction
- LDS
- Leading Decision Summarization
- CE
- Citation Extraction
- SBERT
- Sentence-Bert
- Doc2Doc
- Document-to-Document
- mUSE
- Multilingual Universal Sentence Encoder
1 Introduction
Neural Machine Translation (NMT) is one of the most studied Natural Language Processing (NLP) tasks. From encoder-decoder pipelines (Dai and Le, 2015; Vaswani et al., 2017) to modern decoder-only models (Brown et al., 2020; Touvron et al., 2023) NMT systems based on large language models (LLMs) have in recent years achieved notable advancements in translating texts across various genres (Ou et al., 2023; Zhang et al., 2024; Han et al., 2024) and in both high- and low-resource languages (Moslem et al., 2023; Vilar et al., 2023; Alves et al., 2023; Oliver et al., 2024). Nevertheless, the shortage of high-quality multilingual parallel legal translation data for training LLMs has hindered the performance of state-of-the-art NMT systems in translating legal texts. This limitation is primarily due to the discourse structures (Wiesmann, 2019) and specialized terminology (Katz et al., 2023) of legal texts, which consequently result in the current limited degree of automation for translation in the legal domain.
In multilingual countries like Switzerland, where legal documents are primarily translated manually by experts, developing reliable NMT systems for legal texts would significantly improve governmental efficiency and reduce administrative bottlenecks (Martínez-Domínguez et al., 2020). Beyond operational benefits, such systems could democratize access to legal information by enabling faster and more cost-effective translations across multiple national languages. Especially in lower-resourced languages like Romansh where full translation coverage is not currently economical, support from high-quality NMT systems could be game-changing. This broader accessibility would enhance the transparency of political decision-making and promote more inclusive civic participation (Moniz and Escartín, 2023). The potential impact extends beyond government operations to the private sector where law firms and businesses operating across linguistic regions could benefit from improved legal translation capabilities, potentially reducing costs and accelerating legal processes while maintaining accuracy. Although initial efforts have been made to develop NMT systems for translating Swiss legal documents (Martínez-Domínguez et al., 2020; Canavese and Cadwell, 2024), it remains unclear how well current LLMs perform on large benchmarks for translating Swiss legal texts, both in zero-shot and fine-tuning settings.
To address the shortage of Swiss legal training data and advance legal translation, we present three main contributions:
1.
SwiLTra-Bench: A large-scale benchmark of over 180K aligned Swiss legal translation pairs (laws, court decisions, press releases) spanning five languages (the four official Swiss languages plus English), substantially expanding available training data.
2.
Comprehensive Model Comparison: The first large-scale evaluation of frontier LLMs and fine-tuned open SLMs on Swiss legal translations in both zero-shot and fine-tuning settings, providing insights into their relative strengths.
3.
SwiLTra-Judge: An LLM-based method aligned with human expert annotations, offering a reliable automated framework to assess translation quality.
Our main findings are: a) frontier models consistently perform well across translation tasks; b) translation specific systems like MADLAD-400 are strong on laws but fall behind on headnotes; c) fine-tuning open SLMs drastically improves their translation quality but they are still behind zero-shot prompted frontier models; d) translation quality is uniform across languages; and e) agreement among human experts is higher for law translations than for headnotes.
2 SwiLTra-Bench
To support research in NMT, text alignment, and legal document processing, we present SwiLTra-Bench—a dataset uniting original legal texts and press releases on key court rulings.
| Source | Split | #file | #de | #fr | #it | #rm | #en |
| Law | Train | 5,206 | 5,206 | 5,206 | 5,206 | 51 | 219 |
| Valid | 10 | 10 | 10 | 10 | 10 | 10 | |
| Test | 20 | 20 | 20 | 20 | 20 | 20 | |
| Article | Train | 129,070 | 126,308 | 127,049 | 126,223 | 8,680 | 16,347 |
| Valid | 789 | 785 | 785 | 784 | 785 | 785 | |
| Test | 740 | 738 | 738 | 738 | 738 | 738 | |
| Paragraph | Train | 153,970 | 145,106 | 146,953 | 145,267 | 19,556 | 32,499 |
| Valid | 1,490 | 1,441 | 1,438 | 1,437 | 1,441 | 1,439 | |
| Test | 1,214 | 1,176 | 1,178 | 1,178 | 1,177 | 1,176 |
| Source | Split | #file | #de | #fr | #it |
| BGE | Train | 13,330 | 13,330 | 13,330 | 13,330 |
| Valid | 1,900 | 1,900 | 1,900 | 1,900 | |
| Test | 3,801 | 3,801 | 3,801 | 3,801 | |
| Regest | Train | 13,550 | 13,550 | 13,550 | 13,550 |
| Valid | 1,924 | 1,924 | 1,924 | 1,924 | |
| Test | 3,890 | 3,890 | 3,890 | 3,890 | |
| Text | Train | 26,008 | 26,008 | 26,008 | 26,008 |
| Valid | 3,805 | 3,805 | 3,805 | 3,805 | |
| Test | 7,316 | 7,316 | 7,316 | 7,316 |
| Source | Split | #file | #de | #fr | #it |
| Press | Train | 867 | 867 | 867 | 152 |
| Valid | 100 | 100 | 100 | 100 | |
| Test | 200 | 200 | 200 | 200 |
2.1 Data Collection
SwiLTra-Bench contains three sub-datasets:
1.
Swiss Law Translations (CH-Law-Trans), including law-level (entire legal documents), article-level (individual articles), and paragraph-level (paragraphs within articles) translations.
2.
Headnote Translations (CH-Headnote-Trans) of Swiss Supreme Court landmark court decisions (“Bundesgerichtsentscheide” (BGE) in German, “Arrêts du Tribunal fédéral” (ATF) in French, and “Decisioni principali del Tribunale federale svizzero” (DFT) in Italian) at the BGE-level (complete summaries of court decisions), regest-level (summaries focused on core legal issues), and text-level (detailed extraction of specific legal statements).
3.
Swiss Supreme Court Press Release Translations (CH-Press-Trans).
All datasets contain parallel translations in German (de), French (fr), and Italian (it). Additionally, for CH-Law-Trans, some documents contain translations in Romansh (rm) and English (en).
We provide details of the data structure with concrete dataset examples in Appendix D.
2.2 Dataset Splits
We first segment each dataset by a unique identifier (entire laws, entire headnotes, and entire press releases) to ensure that no single law, headnote, or press release is split across training, validation, and test. For laws, we prioritize examples for the validation and test splits that (1) have more language versions (to guarantee good multilingual coverage), (2) have an official abbreviation (since abbreviations are only set for those laws that are presumed to be cited frequently111https://www.bk.admin.ch/apps/gtr/de/index.html, which we consider a good proxy for practical importance, (3) have shorter text lengths to make evaluation faster and cheaper, and (4) have newer applicability dates so that more recent and multilingual laws are prioritized for validation and testing, resulting in a more realistic evaluation setting. For headnotes, we similarly prioritize those with more recent publication years for validation and test. Finally, for press releases, we focus on maximizing multilingual coverage by ensuring all validation and test examples are available in all present languages (German, French and Italian). The training sets contain all examples not held out for validation or testing.
2.3 Data Statistics
Table 1 presents the overall statistics of the three datasets included in SwiLTRa-Bench. We visualize the training set text lengths for the shortest levels used for training and evaluation in Figure B.1. For completeness, we show histograms for all levels in Figure B.2 and Figure B.3. To calculate these statistics, we used an NLTK222https://www.nltk.org word tokenizer, splitting sentences based on whitespace and punctuation.
3 Experimental Setup
3.1 Evaluation
To paint a representative picture of translation capabilities, we evaluate models across five main classes: 1) translation models, i.e., models specifically trained for translation tasks, 2) frontier models, i.e., large foundation models pre-trained on web-scale data and post-trained on diverse tasks, 3) reasoning models, i.e., models using significant resources at test time to improve output quality, 4) open models, i.e., typically small language models (SLMs) with publicly available weights, and 5) fine-tuned models, i.e., models specifically fine-tuned on SwiLTRa-Bench. We conducted our evaluation using the lighteval framework due to its ease of use and good support for custom metrics.333https://github.com/huggingface/lighteval
3.1.1 Metrics
We evaluated translations using lexical (BLEU Papineni et al. (2002), ChrF Popović (2015), METEOR Banerjee and Lavie (2005)) and model-based metrics (BERTScore Zhang et al. (2020), BLEURT Sellam et al. (2020), XCOMET Guerreiro et al. (2024), GEMBA-MQM Kocmi and Federmann (2023)). Due to the 512-token limit, BLEURT and XCOMET cannot process press releases. Given GEMBA-MQM’s strong correlation with human judgments, we prioritized it alongside XCOMET, METEOR, and ChrF, ensuring both lexical and trained metrics for diversity.
3.2 Fine-tuning
To provide an overview of the current open SLM landscape, we fine-tuned Gemma-2 2B and 9B Team et al. (2024), Llama 1B, 3B and 8B Grattafiori et al. (2024), Phi-3.5 mini and Phi-3 medium Abdin et al. (2024), and Qwen2.5 0.5B, 1.5B, 3B, 7B, 14B and 32B Team (2024) models on our dataset. We fine-tuned using Hugging Face transformers444https://github.com/huggingface/transformers and unsloth555https://github.com/unslothai/unsloth using 4-bit quantization and 8bit AdamW Loshchilov and Hutter (2019); Dettmers et al. (2022) on a single 80GB NVIDIA H100 GPU. We used rank stabilized LoRA Hu et al. (2021); Kalajdzievski (2023) with rank 16 and alpha 16. We trained with the model’s native chat template on sequence length 512, covering more than 99% of the training dataset and truncating the rest. The instruction template was simply:
{source language}: {source text} {target language}: {target text}
We trained on the entire training set for maximal coverage of the data. We used packing, weight decay 0.01, batch size 128 and early stopping with patience 3. In most cases, the lowest evaluation loss is reached after exactly 1 epoch. We used a linear learning rate schedule with 1000 warmup steps and learning rate . We manually tuned the learning rate ( - ), weight decay (0.01, 0.1), label smoothing (factor 0, 0.01, 0.1) and LoRA rank (16, 128). We used the train and validation sets of the Law and Headnote translations on the lowest (shortest) levels, i.e., the paragraph and text levels. For all fine-tuned models, we used the instruction-tuned variant since they have shown to better adapt to new tasks Niklaus et al. (2024).
4 Results and Analysis
In the tables, we bolded the highest and underlined the second highest score per metric. Unless stated otherwise, we excluded Romansh from the evaluations to ensure comparability, since it is not supported by the translation models. Unless stated otherwise, results are averaged over source languages, target languages, and tasks. In general, we considered the law and headnote translation tasks at the highest granularity (paragraph-level and text-level) so we can compare all model categories (translation models and fine-tuned models are optimized for shorter sequence lengths). We show all metrics with standard errors obtained through bootstrapping. Higher values are better for all metrics.
4.1 Statistical Significance Testing
The model comparisons presented throughout this section are descriptive in nature. When we state that one model “outperforms” or “outcompetes” another, we are describing observed differences in the reported metrics; we do not claim statistical significance unless explicitly stated. We only apply null hypothesis significance testing (NHST) in those instances where we explicitly report statistically significant results. In such cases, we provide information about the tests used, statistics obtained, and effect sizes.
4.2 Translation Models
We compare translation models in Table 2. Surprisingly, Google-Translate performs poorly compared to open translation models like MADLAD-400 Kudugunta et al. (2024) and Tower-Instruct. Facebook’s SeamlessM4T Communication et al. (2023) model’s text-to-text capabilities also underwhelm. MADLAD-400 performs very well, outperforming GPT-4o on XCOMET. The Tower Alves et al. (2024) models land somewhere in between.
| Model | Size | ↑ GEMBA-MQM | ↑ XCOMET | ↑ METEOR | ↑ ChrF |
| Google-Translate | N/A | 53.20 ± 0.2 | 64.61 ± 0.1 | 41.15 ± 0.1 | 47.81 ± 0.1 |
| MADLAD-400-3B | 3B | 62.89 ± 0.1 | 86.82 ± 0.1 | 42.44 ± 0.1 | 51.36 ± 0.1 |
| MADLAD-400-7B | 7B | 62.66 ± 0.1 | 87.40 ± 0.1 | 43.70 ± 0.1 | 51.67 ± 0.1 |
| MADLAD-400-10B | 10B | 61.46 ± 0.1 | 86.65 ± 0.1 | 43.10 ± 0.1 | 52.24 ± 0.1 |
| SeamlessM4T | 2B | 23.35 ± 0.2 | 43.03 ± 0.1 | 37.81 ± 0.1 | 24.90 ± 0.1 |
| TowerInstruct-7B | 7B | 54.04 ± 0.2 | 72.97 ± 0.1 | 41.65 ± 0.2 | 43.00 ± 0.1 |
| TowerInstruct-13B | 13B | 57.38 ± 0.2 | 75.94 ± 0.1 | 43.95 ± 0.2 | 48.46 ± 0.1 |
4.3 Frontier Models
We show results for frontier and reasoning models in Table 3. GPT-4o underperforms both of its peers Claude-3.5-Sonnet and Llama-3.1-405B. This is particularly unexpected, as models tend to favor their own completions Panickssery et al. (2024), and GEMBA-MQM is operated by GPT-4o. Claude-3.5-Sonnet demonstrates strong performance, competing closely with o1, the top-performing model. Surprisingly, o1-mini performs only on par with the other models at the smaller scale and even underperforms Claude-3.5-Haiku. Overall, Anthropic’s models are really strong, and even more so from a cost-to-performance perspective compared to reasoning models like o1.
| Model | Size | ↑ GEMBA-MQM | ↑ XCOMET | ↑ METEOR | ↑ ChrF |
| Claude-3.5-Sonnet | large | 80.66 ± 0.2 | 90.70 ± 0.1 | 56.71 ± 0.2 | 65.87 ± 0.1 |
| DeepSeek-V3 | large | 80.04 ± 0.2 | 89.77 ± 0.1 | 56.60 ± 0.1 | 69.99 ± 0.1 |
| DeepSeek-R1 | large | 77.90 ± 0.2 | 84.36 ± 0.1 | 55.79 ± 0.1 | 69.12 ± 0.1 |
| GPT-4o | large | 80.27 ± 0.2 | 80.96 ± 0.1 | 55.56 ± 0.1 | 63.27 ± 0.1 |
| Gemini-1.5-Pro | large | 81.88 ± 0.2 | 87.13 ± 0.1 | 57.92 ± 0.1 | 70.07 ± 0.1 |
| Llama-3.1-405B | large | 81.59 ± 0.1 | 89.37 ± 0.1 | 54.48 ± 0.1 | 68.07 ± 0.1 |
| Mistral-Large | large | 81.88 ± 0.2 | 87.04 ± 0.1 | 54.86 ± 0.1 | 63.71 ± 0.1 |
| o1 | large | 85.81 ± 0.1 | 91.35 ± 0.1 | 58.91 ± 0.1 | 70.11 ± 0.1 |
| Claude-3.5-Haiku | small | 80.40 ± 0.2 | 88.84 ± 0.1 | 52.15 ± 0.2 | 61.09 ± 0.1 |
| GPT-4o-mini | small | 82.59 ± 0.2 | 87.90 ± 0.1 | 54.03 ± 0.1 | 59.86 ± 0.1 |
| Gemini-1.5-Flash | small | 80.76 ± 0.2 | 85.33 ± 0.1 | 55.35 ± 0.1 | 65.44 ± 0.1 |
| Llama-3.3-70B | small | 79.25 ± 0.2 | 88.02 ± 0.1 | 53.43 ± 0.1 | 65.92 ± 0.1 |
| Mistral-Small | small | 81.69 ± 0.2 | 87.04 ± 0.1 | 54.83 ± 0.1 | 63.66 ± 0.1 |
| o1-mini | small | 81.96 ± 0.2 | 87.46 ± 0.1 | 53.34 ± 0.1 | 59.32 ± 0.1 |
4.4 Fine-tuned Models
Fine-tuning leads to notable performance gains (see Appendix Table C.1). Figure 2 presents fine-tuned models’ performance across various sizes. The two Gemma models and particularly the Llama 1B and 3B models, advance the Pareto frontier, though performance starts to flatten at the 3B scale and plateaus after 9B parameters. Interestingly, both Phi models clearly underperform their peers.
4.5 Performance Progression by Model Size
The Qwen2.5 model family, with six sizes from 0.5B to 32B parameters, is ideal for studying performance progression over model size. We analyzed fine-tuned Qwen models up to 32B using five metrics (two model-based, three lexical) in Figure 3. METEOR is the only lexical metric well correlated with XCOMET and GEMBA-MQM. All three confirm a clear trend that larger models produce higher-quality translations. GEMBA-MQM shows the largest score range (GEMBA-MQM: 52.4 - 82.8 vs XCOMET: 69.5 - 87.9 and METEOR: 56.8 - 65.1) and making it most useful for differentiating models. Interestingly, both ChrF and BLEU are negatively correlated with the model-based metrics on this task for the fine-tuned Qwen models. Beyond the inherent subjectivity in assessing translation quality, this may hint at the greater importance of a legal text’s conveyed meaning over the mere use of certain exact terms.
4.6 Comparison Across Tasks
In Table 4, we show the best models’ performance per category across tasks. The best open small model falls far behind the others but, with fine-tuning, overtakes the best translation model. It matches the smaller frontier models but still lags behind the larger ones. All models except MADLAD-400-7B perform better on headnote than law translation. While Sonnet competes with o1 on headnote and law translation, it falls off on press releases.
| Model | Category | Task | ↑ GEMBA-MQM | ↑ METEOR | ↑ ChrF |
| o1 | reasoning | Headnote | 93.50 ± 0.1 | 60.89 ± 0.1 | 62.62 ± 0.2 |
| o1 | reasoning | Law | 91.11 ± 0.1 | 55.87 ± 0.1 | 66.84 ± 0.1 |
| o1 | reasoning | Press | 64.62 ± 0.4 | 59.28 ± 0.3 | 78.38 ± 0.1 |
| Claude-3.5-Sonnet | frontier | Headnote | 88.65 ± 0.1 | 61.39 ± 0.1 | 63.96 ± 0.2 |
| Claude-3.5-Sonnet | frontier | Law | 85.71 ± 0.1 | 52.16 ± 0.1 | 73.15 ± 0.1 |
| Claude-3.5-Sonnet | frontier | Press | 60.83 ± 0.8 | 55.29 ± 0.5 | 55.47 ± 0.1 |
| MADLAD-400-7B | translation | Headnote | 80.54 ± 0.2 | 57.71 ± 0.1 | 67.49 ± 0.2 |
| MADLAD-400-7B | translation | Law | 85.06 ± 0.2 | 57.09 ± 0.2 | 61.86 ± 0.2 |
| SLT-Qwen2.5-32B | finetuned | Headnote | 82.58 ± 0.1 | 66.56 ± 0.1 | 75.17 ± 0.2 |
| SLT-Qwen2.5-32B | finetuned | Law | 80.80 ± 0.2 | 64.41 ± 0.1 | 76.90 ± 0.1 |
| Qwen2.5-14B | open | Headnote | 69.88 ± 0.2 | 47.09 ± 0.1 | 53.58 ± 0.2 |
| Qwen2.5-14B | open | Law | 63.04 ± 0.2 | 34.33 ± 0.1 | 52.02 ± 0.1 |
4.7 Comparison Across Languages
In Figure 4, we compare the best models per category across language directions on CH-Law-Trans. Performance to and from German, French, and Italian is homogeneous across models. When translating from English to the other languages, all models perform worse than from the three main Swiss languages. Since the English source texts are already translations and are not legally binding, the federal translators may have applied less rigor in generating them, potentially resulting in lower quality and slight deviations. Anecdotally, the lawyers co-authoring this work confirm that the English source texts are occasionally less precise. So, the lower scores may also indicate that the judge model bases its grading on imperfect source text.
Romansh is a low-resource language and only spoken by less than 50K people in Switzerland.666https://www.bfs.admin.ch/bfs/de/home/statistiken/bevoelkerung/sprachen-religionen/sprachen.html As our dataset consists of the entire data readily available for federal laws, Supreme Court headnotes, and Supreme Court press releases in Switzerland, it is not possible to extend the Romansh coverage. Adding additional data sources is not a viable option due to likely lower data quality and limited sentence level alignment across languages. Romansh is not supported by most translation models such as MADLAD-400. In our opinion, the fact that these models don’t support Romansh at the moment highlights the value of our dataset for low-resource languages. Indeed, our dataset now enables teams building translation models to train and evaluate on >160K and >8K Romansh translation pairs respectively. Surprisingly, o1 and Sonnet still perform very well when translating from Romansh to other languages. When translating to Romansh, all models’ quality drops off, sometimes sharply. Perhaps similarly to humans, also for LLMs speaking or writing a language seems harder than understanding it.
5 Expert Evaluation
To study how well human legal experts agree with the automated metrics, we conducted an expert evaluation. All experts are authors of the paper; the majority are doctoral candidates, and all hold at least a Bachelor’s degree in Swiss law. We only evaluated the laws and headnotes since they are much shorter and we could evaluate more examples in the time available. Due to limited expert time, we selected the top model from four categories: frontier (Claude-3.5-Sonnet), reasoning (o1), translation (MADLAD-400-7B) and finetuned (SLT-Qwen2.5-32B).
The experts were asked to assign a score between 0 and 100 to each translation. For this purpose, the experts were given a source text, its “gold translation” (official translation of the Swiss authorities) as a reference and a predicted translation. The scores only reflected the completeness and accuracy of the predicted translation, with less emphasis on readability and other stylistic attributes. To ensure consistency, the experts agreed on a point deduction system in advance and discussed certain borderline cases (annotation guidelines are in Appendix G). In total, 200 translations were assigned a score by exactly two experts. Each expert assigned scores independently, without consulting the other annotators. For the expert agreement with judge metrics (see section 6) and for the evaluation of the best models (see subsection 5.2), we averaged the scores of the two annotators.
5.1 Inter-Annotator Agreement
The average Krippendorff’s was 0.56 for laws and 0.41 for headnotes. Agreement was generally higher for laws than for headnotes across most language pairs and tasks. This consistent pattern is unlikely due to a calibration issue and instead suggests that the headnote task inherently allows for more subjectivity. This is likely due to headnotes being longer on average and containing greater interpretive nuance, which can lead annotators to miss or accord different importance to different details. We also hypothesize that laws inherently exhibit a more structured and clearer pattern, making them easier to translate and evaluate.
The moderate inter-annotator agreement suggests that, despite clear instructions, a certain degree of subjectivity was inherent in the task. In addition, we observed smaller differences between the individual language pairs, suggesting that not all annotators were perfectly aligned. However, disagreements tended to be minor and were rarely fundamental. In Figure 5 we show the absolute point difference between the two annotators evaluating the same samples. In almost half of the cases the two annotators completely agree and in 92% the difference is smaller than 30 points.
5.2 Which Model is the Best?
In Table 5 we show the expert scores together with the best metrics for the best models per category. It is evident here that XCOMET aligns best with the experts. We conclude that both for translating laws and headnotes Claude 3.5 Sonnet is the best model followed by o1 for laws and both o1 and the finetuned Qwen2.5-32B model for headnotes.
| Model | Task | ↑ Experts | ↑ XCOMET | ↑ BLEURT | ↑ GEMBA-MQM |
| Claude-3.5-Sonnet | Headnote | 89.21 ± 2.2 | 90.91 ± 1.5 | 28.96 ± 3.7 | 86.53 ± 1.6 |
| Claude-3.5-Sonnet | Law | 94.55 ± 1.1 | 93.30 ± 1.1 | 34.16 ± 3.2 | 88.86 ± 1.2 |
| MADLAD-400-7B | Headnote | 71.77 ± 3.3 | 85.57 ± 2.8 | 12.20 ± 3.1 | 76.13 ± 4.9 |
| MADLAD-400-7B | Law | 83.77 ± 2.8 | 89.42 ± 2.3 | 28.97 ± 3.5 | 88.63 ± 2.0 |
| SLT-Qwen2.5-32B | Headnote | 84.86 ± 2.4 | 88.62 ± 2.1 | 30.78 ± 4.3 | 75.89 ± 4.2 |
| SLT-Qwen2.5-32B | Law | 85.74 ± 2.1 | 88.03 ± 2.2 | 36.42 ± 3.9 | 81.78 ± 2.5 |
| o1 | Headnote | 84.29 ± 2.1 | 89.58 ± 1.8 | 16.77 ± 4.2 | 92.34 ± 1.4 |
| o1 | Law | 89.91 ± 1.5 | 92.33 ± 1.5 | 28.19 ± 3.2 | 92.97 ± 1.0 |
6 SwiLTra-Judge
Automatic evaluation of natural language generation is challenging. Lexical metrics like BLEU or METEOR correlate weakly with human judgments Zhang et al. (2020). Early model-based metrics such as BERTScore or BLEURT perform better, but recently, LLM-as-Judge has emerged as the dominant paradigm Zheng et al. (2023). Each task, however, is unique and requires its own judge setup. In this section, we ablate key aspects of the judge setup, including the judge model, prompt, and few-shot sample selection.
6.1 Setup
We use GPT-4o, GPT-4o-mini, Gemini-1.5-pro, and Gemini-1.5-flash in our judge model ablation. We also tested Claude Sonnet and Haiku as judges, but they failed to follow grading instructions.777They would insist on generating JSON output while we very clearly just asked for plain-text. The o1 and o1-mini models showed very low or even negative correlations with human judgments and are thus excluded. We randomly selected one few-shot example from the dev sets of laws, headnotes, and press releases. To ensure judge models saw diverse translation qualities, we chose models of varying strengths (Claude 3.5 Sonnet for laws, Mixtral-8x7B-Instruct-v0.1 for headnotes, and Qwen2.5-1.5B-Instruct for press releases). We used a simple prompt “Translate to target-language” to generate translations. Sample judgments per few-shot example were written by one lawyer author and double-checked by another. We tested two few-shot styles single (all examples in one language direction: fr-de) and diverse (law article en-it, headnote de-fr and press release fr-de). We ablated two user prompts with absolute grading (basic and detailed) and one with deduction grading similar to the codebook given to the human expert annotators (codebook). Judge prompts are in Appendix F.
We measured the correlation of our judge setups with the human expert scores on the 400 human annotated samples. To get a higher confidence signal, we removed samples where the two human experts disagreed by 30 points or more (32/400 or 8%). Find complete results in Table E.1. Unless specified otherwise, we report Spearman correlation with human judgments with cross validation. Based on our expert evaluation, we answer the following research questions (RQs):
RQ1: Are small models judges good enough?
A: Yes, the small models even outperform their larger counterparts. Over all tested configurations GPT-4o and GPT-4o-mini are tied at mean Spearman correlation. Gemini-1.5-flash even outperforms Gemini-1.5-pro as a judge model ( vs ). For the best configuration GPT-4o-mini even outperforms GPT-4o ( vs ) and the same holds for Gemini-1.5-flash vs Gemini-1.5-pro ( vs ).
RQ2: Is the deduction judgment style better than the absolute style?
A: Judges using the deduction style align more closely with human judgments. Across all configurations, there is little difference between the two absolute styles ( for basic and for detailed user prompt). However, the deduction style aligns much more closely with experts (). The top six highest correlating configurations all use the deduction style. This finding anecdotally confirms that LLM judge models reach judgments more similar to human experts when prompted in a more aligned way.
RQ3: Are few-shot examples in a single language pair sufficient, or is it necessary to include examples from diverse language pairs?
A: On average, the language directions of the few shot examples do not matter, but the best configuration uses diverse language directions. Across all 24 investigated configurations, there is only minimal difference between the single and diverse language direction setup ( vs ). However, the best configuration overall, uses diverse language directions.
RQ4: How does SwiLTra-Judge perform compared to other metrics?
A: Our SwiLTra-Judge exhibits the highest correlation with human judgments among tested translation metrics. Figure 6 shows Spearman correlation with human judgments for sample-level metrics (this excludes BLEU and ChrF). As expected, METEOR and BERTScore perform poorly, with correlations below 0.2. Surprisingly, the recent GEMBA-MQM metric both underperforms BLEURT and XCOMET. Our SwiLTra-Judge-Single (gemini-1-5-flash-codebook-diverse-deduction) is significantly better than the second-best metric XCOMET ( vs , ). Our SwiLTra-Judge-Ensemble (a simple mean score from GPT-4o-mini and Gemini 1.5 Flash with codebook style and both single and diverse few-shot setup) is again significantly better than our SwiLTra-Judge-Single ( vs , ). To compute both p-values, we have performed an unpaired t-test with each.
6.2 Judge Harshness
In Figure 7 we show the average score over the best four generator models per category across judge models, system and few shot styles. We confirm here that the language directions of the few shot examples only has a minor effect. We see that the detailed system style leads to the harshest scores across models. Interestingly, Gemini-1.5-pro and GPT-4o-mini judge very similarly in terms of harshness. All models except GPT-4o judge more leniently with the codebook system style.
6.3 Best Model Per Task
Now that we built a trusted metric for our translation benchmark, we ran it over the entire dataset for all models. With this, we can recommend the best model for each task. In Figure 1 we show the top three models per task using SwiLTra-Judge as a metric. There are no large differences among top models, but the highest scores are achieved by Sonnet on headnotes, MADLAD-400-7B on laws and o1 on press releases. Sonnet ranks in the top 3 for all tasks. One reason for Sonnet’s high scores could be that the few-shot example in the judge prompt with the highest score was translated by Sonnet, possibly making the judge models prefer its style. However, human experts clearly favored Sonnet without bias from few-shot examples.
7 Related Work
The application of NLP to legal texts has seen significant growth in recent years. This increased attention is driven by the growing need to automate and enhance legal processes, improve access to justice, and handle the vast amounts of legal documentation produced globally.
Recent research has explored various aspects of legal text processing. Legal judgment prediction has emerged as a crucial area, with studies demonstrating success across different jurisdictions, including the US Semo et al. (2022), Europe Vaudaux et al. (2023) and Switzerland Niklaus et al. (2021, 2022). Notable advances have been made in verdict prediction Medvedeva et al. (2020), topic classification Papaloukas et al. (2021); Benedetto et al. (2023); Rasiah et al. (2023), and legal QA systems Zhong et al. (2020). LegalBench Guha et al. (2023) LexGLUE Chalkidis et al. (2022) and LEXTREME Niklaus et al. (2023) are established as comprehensive benchmark suites comprising multiple legal NLP tasks, including text classification, named entity recognition, and legal entailment across various jurisdictions and legal areas.
The translation of legal texts has significant societal impact and is increasingly important for training translators and practical applications, especially as machine translation gains prominence Killman (2024). However, legal translation poses challenges due to domain-specific terminology, reliability in legal formulae, and non-compliance with legal conventions Killman (2023); Giampieri (2023). While some U.S. courts have considered NMT, it remains far from replacing human translators Vieira et al. (2021). Robust NMT systems are essential for judicial and governmental services, with recent advancements leveraging pretrained LLMs and fine-tuning techniques Zhu et al. (2024). Prior research has focused on legal NMT for languages like Chinese, and Arabic Ding (2024); ElFqih and Monti (2023). However, a significant research gap remains in translating legal texts between Switzerland’s national languages, which our work aims to address.
8 Conclusions
In this work, we introduced SwiLTra-Bench, a high-quality multilingual legal translation benchmark, and evaluated mainstream LLM-based NMT systems under both zero-shot and fine-tuned settings. Our analysis, validated by human expert annotations, showed that frontier models outperform all others, while translation-specific systems like MADLAD-400 excel on laws but struggle with headnotes. Fine-tuning open LLMs significantly improves their performance, though they still lag behind zero-shot frontier models, and translation quality remains consistent across Swiss languages. Finally, our SwiLTra-Judge model, optimized for legal translation evaluation, achieves the highest alignment with human expert judgments, providing a valuable automated metric for future research.
Limitations & Future Work
Our fine-tuned models are much stronger than the initial instruction-tuned open models they are based on, but they still under-perform large closed models. Future work could investigate techniques such as model merging Yang et al. (2024) to further improve and bring them closer to the frontier models. While we evaluated a large variety of models, we could not evaluate them all. Future work could investigate other promising models such as Grok.888https://x.ai/blog/grok-2 We took great care to validate our results with human expert studies. However, our resources were limited and we could not investigate certain languages (e.g., Romansh) and our sample sizes were still rather small. Future work could perform a more broad and in-depth human evaluation.
Ethics Statement
Our benchmark contains no personal, sensitive, or private information; it consists solely of publicly available data.
Acknowledgments
We thank the anonymous reviewers for their constructive feedback. Luka Nenadic’s research is funded by SNSF grant No. 10002634.
Appendix A Use of AI Assistants
We used GPT-4o and Claude Sonnet 3.5 for coding, shortening texts and editing LaTeX more efficiently.
Appendix B Corpus Distribution of Text Lengths
Appendix C Additional Experimental Results
| Model | Family | Category | Size | ↑ GEMBA-MQM | ↑ XCOMET | ↑ BLEURT | ↑ METEOR | ↑ ChrF |
| Gemma-2-2B | Gemma | open | 2B | 9.90 ± 0.1 | 35.52 ± 0.1 | -102.08 ± 0.3 | 6.97 ± 0.1 | 11.12 ± 0.1 |
| SLT-Gemma-2-2B | Gemma | finetuned | 2B | 72.71 ± 0.2 | 82.39 ± 0.1 | 26.72 ± 0.3 | 61.52 ± 0.1 | 79.48 ± 0.1 |
| Gemma-2-9B | Gemma | open | 9B | 12.15 ± 0.1 | 36.54 ± 0.1 | -102.19 ± 0.3 | 7.48 ± 0.1 | 0.00 ± 0.1 |
| SLT-Gemma-2-9B | Gemma | finetuned | 9B | 82.54 ± 0.1 | 87.62 ± 0.1 | 32.89 ± 0.2 | 65.16 ± 0.1 | 78.95 ± 0.1 |
| Llama-3.2-1B | Llama | open | 1B | 27.23 ± 0.2 | 48.43 ± 0.2 | -15.64 ± 0.2 | 29.70 ± 0.1 | 39.87 ± 0.1 |
| SLT-Llama-3.2-1B | Llama | finetuned | 1B | 64.14 ± 0.2 | 76.40 ± 0.1 | 26.35 ± 0.2 | 59.03 ± 0.1 | 79.76 ± 0.1 |
| Llama-3.2-3B | Llama | open | 3B | 54.13 ± 0.2 | 67.43 ± 0.2 | 0.59 ± 0.2 | 38.57 ± 0.1 | 50.47 ± 0.1 |
| SLT-Llama-3.2-3B | Llama | finetuned | 3B | 75.56 ± 0.2 | 83.39 ± 0.1 | 30.47 ± 0.2 | 62.54 ± 0.1 | 79.32 ± 0.1 |
| Llama-3.1-8B | Llama | open | 8B | 67.09 ± 0.2 | 75.03 ± 0.2 | 6.25 ± 0.2 | 43.72 ± 0.1 | 50.56 ± 0.1 |
| SLT-Llama-3.1-8B | Llama | finetuned | 8B | 80.23 ± 0.1 | 86.04 ± 0.1 | 31.91 ± 0.2 | 64.17 ± 0.1 | 80.89 ± 0.1 |
| Phi-3.5-mini | Phi | open | 3.8B | 17.96 ± 0.2 | 41.93 ± 0.1 | -92.40 ± 0.3 | 9.66 ± 0.1 | 11.42 ± 0.1 |
| SLT-Phi-3.5-mini | Phi | finetuned | 3.8B | 73.90 ± 0.2 | 80.31 ± 0.1 | 10.33 ± 0.2 | 56.75 ± 0.1 | 76.72 ± 0.1 |
| Phi-3-medium | Phi | open | 14B | 21.33 ± 0.2 | 38.91 ± 0.1 | -81.04 ± 0.3 | 13.45 ± 0.1 | 17.74 ± 0.1 |
| SLT-Phi-3-medium | Phi | finetuned | 14B | 81.56 ± 0.1 | 87.38 ± 0.1 | 32.39 ± 0.2 | 64.16 ± 0.1 | 80.40 ± 0.1 |
| Qwen2.5-0.5B | Qwen | open | 0.5B | 9.82 ± 0.2 | 41.36 ± 0.2 | -61.96 ± 0.2 | 14.53 ± 0.1 | 28.21 ± 0.1 |
| SLT-Qwen2.5-0.5B | Qwen | finetuned | 0.5B | 52.37 ± 0.2 | 69.48 ± 0.2 | 22.67 ± 0.2 | 56.77 ± 0.1 | 78.66 ± 0.1 |
| Qwen2.5-1.5B | Qwen | open | 1.5B | 35.21 ± 0.2 | 58.12 ± 0.2 | -46.89 ± 0.3 | 22.55 ± 0.1 | 36.26 ± 0.1 |
| SLT-Qwen2.5-1.5B | Qwen | finetuned | 1.5B | 69.58 ± 0.2 | 80.43 ± 0.1 | 26.90 ± 0.2 | 60.45 ± 0.1 | 78.13 ± 0.1 |
| Qwen2.5-3B | Qwen | open | 3B | 48.85 ± 0.2 | 61.18 ± 0.2 | -11.88 ± 0.2 | 34.77 ± 0.1 | 44.46 ± 0.1 |
| SLT-Qwen2.5-3B | Qwen | finetuned | 3B | 74.33 ± 0.2 | 82.99 ± 0.1 | 27.78 ± 0.2 | 61.61 ± 0.1 | 78.03 ± 0.1 |
| Qwen2.5-7B | Qwen | open | 7B | 58.79 ± 0.2 | 69.07 ± 0.2 | 0.73 ± 0.2 | 39.41 ± 0.1 | 42.67 ± 0.1 |
| SLT-Qwen2.5-7B | Qwen | finetuned | 7B | 78.96 ± 0.1 | 86.03 ± 0.1 | 31.07 ± 0.2 | 63.40 ± 0.1 | 77.54 ± 0.1 |
| Qwen2.5-14B | Qwen | open | 14B | 72.70 ± 0.2 | 79.54 ± 0.1 | 9.27 ± 0.2 | 45.06 ± 0.1 | 56.13 ± 0.1 |
| SLT-Qwen2.5-14B | Qwen | finetuned | 14B | 82.78 ± 0.1 | 87.46 ± 0.1 | 32.37 ± 0.2 | 64.73 ± 0.1 | 78.56 ± 0.1 |
| Qwen2.5-32B | Qwen | open | 32B | 70.30 ± 0.2 | 76.34 ± 0.1 | 6.91 ± 0.2 | 45.33 ± 0.1 | 57.94 ± 0.1 |
| SLT-Qwen2.5-32B | Qwen | finetuned | 32B | 82.77 ± 0.1 | 87.90 ± 0.1 | 33.20 ± 0.2 | 65.07 ± 0.1 | 76.75 ± 0.1 |
Appendix D Corpus Examples
| Dataset | Field | Comment |
| CH-Press-Trans | filename | Unique identifier of each press release |
| de_text | Press release content in German | |
| fr_text | Press release content in French | |
| it_text | Press release content in Italian | |
| has_all_langs | Binary indicator of language availablity | |
| CH-Law-Trans | abbreviation | The abbreviation of the law |
| url | URL linking to the legal text on Fedlex | |
| rsNr | Swiss federal register number | |
| artNr | Article number | |
| parNr | Paragraph number | |
| dateApplicability | Date of applicability of the law | |
| {de/fr/it/rm/en}_lawTitle | Law titles in different languages | |
| {de/fr/it/rm/en}_artTitle | Article titles in different languages | |
| {de/fr/it/rm/en}_lawText | Full law texts in different languages | |
| {de/fr/it/rm/en}_artText | Full article texts in different languages | |
| {de/fr/it/rm/en}_parText | Full paragraph texts in different languages | |
| {de/fr/it/rm/en}_lawHtml | Law texts in HTML format in different languages | |
| {de/fr/it/rm/en}_artHtml | Article texts in HTML format in different languages | |
| {de/fr/it/rm/en}_parHtml | Paragraph texts in HTML format in different languages | |
| CH-Headnote-Trans | bge | Case identifier |
| year | Year of the court decision | |
| volume | Volume number of the court decision | |
| pageNumber | Page number of the court decision | |
| regesteNumber | Number assigned to the regeste | |
| textNumber | Number assigned to the specific text extract | |
| {de/fr/it}_bgeText | Full summary texts in different languages | |
| {de/fr/it}_regesteText | Regeste texts in different languages | |
| {de/fr/it}_regesteTitle | Regeste title in different languages | |
| {de/fr/it}_text | Text extract in different languages |
[!htb]
An Example of CH-Law-Trans: Article Dataset
[!htb]
An Example of CH-Headnote-Trans: BGE Dataset {listing*}[!htb]
An Example of CH-Press-Trans Dataset
Appendix E Judge Correlations
| Metric | Spearman (Bootstrap) | Spearman (CV) | RMSE (CV) | MAE (CV) |
| judge-ensemble | 0.536 [0.453, 0.608] | 0.533 ± 0.080 | 16.090 ± 1.424 | 12.979 ± 0.882 |
| gemini-1-5-flash-codebook-diverse-deduction | 0.506 [0.424, 0.586] | 0.504 ± 0.074 | 15.215 ± 2.216 | 11.100 ± 0.670 |
| XCOMET-XXL | 0.484 [0.403, 0.567] | 0.477 ± 0.093 | 14.877 ± 1.372 | 10.204 ± 0.748 |
| gpt-4o-mini-codebook-single-deduction | 0.474 [0.387, 0.551] | 0.477 ± 0.095 | 22.168 ± 3.064 | 16.944 ± 1.691 |
| gemini-1-5-flash-codebook-single-deduction | 0.461 [0.374, 0.542] | 0.466 ± 0.069 | 16.049 ± 2.212 | 10.990 ± 0.720 |
| gpt-4o-mini-codebook-diverse-deduction | 0.459 [0.378, 0.538] | 0.459 ± 0.094 | 22.527 ± 2.678 | 17.138 ± 1.113 |
| gpt-4o-codebook-single-deduction | 0.444 [0.352, 0.538] | 0.447 ± 0.070 | 29.020 ± 3.982 | 19.606 ± 1.951 |
| gpt-4o-codebook-diverse-deduction | 0.427 [0.334, 0.516] | 0.412 ± 0.044 | 30.902 ± 1.843 | 21.537 ± 0.555 |
| gpt-4o-detailed-single-absolute | 0.418 [0.330, 0.501] | 0.427 ± 0.052 | 35.940 ± 4.207 | 24.286 ± 3.382 |
| gpt-4o-mini-basic-single-absolute | 0.413 [0.320, 0.497] | 0.422 ± 0.067 | 26.736 ± 1.817 | 21.370 ± 2.036 |
| gpt-4o-basic-single-absolute | 0.411 [0.320, 0.499] | 0.411 ± 0.131 | 20.655 ± 2.947 | 14.596 ± 1.976 |
| gpt-4o-detailed-diverse-absolute | 0.378 [0.284, 0.464] | 0.380 ± 0.090 | 34.786 ± 3.711 | 22.995 ± 3.173 |
| gpt-4o-mini-detailed-single-absolute | 0.378 [0.282, 0.471] | 0.384 ± 0.069 | 35.437 ± 1.866 | 30.302 ± 1.718 |
| gpt-4o-basic-diverse-absolute | 0.376 [0.281, 0.467] | 0.383 ± 0.087 | 21.780 ± 2.249 | 13.550 ± 1.382 |
| gpt-4o-mini-detailed-diverse-absolute | 0.358 [0.259, 0.450] | 0.364 ± 0.097 | 36.480 ± 2.365 | 30.828 ± 2.387 |
| bleurt_large | 0.356 [0.261, 0.445] | 0.364 ± 0.147 | 63.110 ± 5.225 | 58.102 ± 5.064 |
| gpt-4o-mini-basic-diverse-absolute | 0.354 [0.263, 0.447] | 0.361 ± 0.048 | 28.363 ± 1.315 | 22.393 ± 1.347 |
| gemini-1-5-pro-basic-single-absolute | 0.306 [0.209, 0.404] | 0.295 ± 0.083 | 36.203 ± 3.618 | 22.993 ± 3.010 |
| gemini-1-5-pro-codebook-diverse-deduction | 0.303 [0.200, 0.397] | 0.298 ± 0.095 | 34.528 ± 3.882 | 20.580 ± 2.477 |
| GEMBA-MQM_gpt-4o | 0.293 [0.189, 0.383] | 0.289 ± 0.093 | 18.331 ± 1.743 | 12.698 ± 0.787 |
| gemini-1-5-pro-codebook-single-deduction | 0.292 [0.195, 0.392] | 0.292 ± 0.074 | 36.718 ± 2.663 | 21.816 ± 2.158 |
| gemini-1-5-flash-detailed-single-absolute | 0.281 [0.190, 0.375] | 0.275 ± 0.049 | 29.067 ± 3.070 | 18.709 ± 1.651 |
| gemini-1-5-flash-basic-single-absolute | 0.270 [0.179, 0.359] | 0.279 ± 0.110 | 33.283 ± 5.239 | 20.412 ± 3.287 |
| gemini-1-5-flash-basic-diverse-absolute | 0.255 [0.157, 0.351] | 0.249 ± 0.069 | 27.649 ± 5.852 | 16.756 ± 3.233 |
| gemini-1-5-pro-basic-diverse-absolute | 0.252 [0.153, 0.351] | 0.250 ± 0.082 | 36.544 ± 3.470 | 22.990 ± 2.684 |
| gemini-1-5-pro-detailed-diverse-absolute | 0.247 [0.147, 0.345] | 0.250 ± 0.097 | 38.316 ± 2.063 | 26.098 ± 1.148 |
| gemini-1-5-pro-detailed-single-absolute | 0.235 [0.137, 0.334] | 0.244 ± 0.079 | 38.618 ± 2.362 | 27.798 ± 1.977 |
| gemini-1-5-flash-detailed-diverse-absolute | 0.231 [0.130, 0.328] | 0.225 ± 0.055 | 30.091 ± 4.891 | 19.671 ± 3.346 |
| BERTScore-F | 0.163 [0.066, 0.268] | 0.170 ± 0.053 | 36.723 ± 1.340 | 31.523 ± 1.772 |
| meteor | 0.160 [0.057, 0.259] | 0.164 ± 0.125 | 34.170 ± 3.444 | 29.270 ± 3.191 |
Appendix F Judge Prompts
[h]
System Prompt
User Prompt
The system and user prompt of the basic judge setup. {listing*}[h]
System Prompt
User Prompt
The system and user prompt of the detailed judge setup. {listing*}[h]
System Prompt
User Prompt
The system and user prompt of the codebook judge setup.
Appendix G Annotation Guidelines
[h]
The annotation guidelines given to the human experts.